OpenStudy (anonymous):

(using taylor polynomials) Calculate ln(2.5) to 7 correct digits. (hint:use ln(e)=1, e is about 2.718)

OpenStudy (anonymous):

expand about \(e\)?

OpenStudy (anonymous):

\(f(3)=1\) is the constant, then \(f'(x)=\frac{1}{x}\) so \(f'(e)=\frac{1}{e}\) we start with \[f(x)=1+\frac{1}{e}(x-e)\]

OpenStudy (anonymous):

thats what i am thinkin x=e, but i dont know if my answer is correct. My prof, dosent believe in giving out the answers. even to check work

OpenStudy (anonymous):

next term would be \[f''(x)=-\frac{1}{x}\] so \[f''(e)=\frac{1}{e^2}\] we get \[1+\frac{1}{e}(x-e)-\frac{1}{2e^2}(x-e)^2+...\]

OpenStudy (anonymous):

if you ask me it is kind of silly since you cannot do this without a calculator anyway. might as well just find \(\ln(2.5)\)

OpenStudy (anonymous):

can i ask a question? why do i need to approximate ln 2.5 if i get a number when i plug it into a calc?

OpenStudy (anonymous):

like i said, this is kind of silly you cannot compute \(\frac{1}{e}\) without a calculator lets try an experiment

OpenStudy (anonymous):

do you know something about graphing the remainder of a taylor poylnomial?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but you would say that is the right approach right? evalauting the taylor poylnomial at e?

OpenStudy (anonymous):

you wouldnt have by any chance taken numerical anaylsis?

OpenStudy (anonymous):

i think we got close i am pretty sure that is what was in mind for this problem

OpenStudy (anonymous):

we would have to carry the expansion out further to get closer

OpenStudy (anonymous):

how many terms do you think?

OpenStudy (anonymous):

not sure, but the derivatives are easy enough, so i will be ok to do a couple more

OpenStudy (anonymous):

cool, thanks man.

OpenStudy (anonymous):

\[\ln(x)\] \[\frac{1}{x}\] \[-\frac{1}{x^2}\] \[\frac{2}{x^3}\] \[-\frac{6}{x^4}\] for the first ones so \[1,\frac{1}{e},-\frac{1}{e^2},\frac{2}{e^3},-\frac{6}{e^4}\] for the successive derivatives

OpenStudy (anonymous):

i think i got it. it looks like it weill take 6 terms

OpenStudy (anonymous):

\[1+\frac{1}{e}(x-e)-\frac{1}{2e^2}(x-e)^2+\frac{1}{3e^3}(x-e)^3-\frac{1}{4e^4}(x-e)^4\] i think is right

OpenStudy (anonymous):

doing that gives us 0.9162907

OpenStudy (anonymous):

comparing that to the ln 2.5 i think that 7 digits do match up

OpenStudy (anonymous):

but whats giving me fits is that i have to use the taylor remainder to prove that i have accuracy up to that 7th digit

OpenStudy (anonymous):

pattern becomes clear you get yeah looks like it .9162907318 according to wolfram

OpenStudy (anonymous):

what is the bound on the remainder?

OpenStudy (anonymous):

isn't it just the next term? or am i mistaken?

OpenStudy (anonymous):

yeah its, k+1

OpenStudy (anonymous):

so if we take it out this far, next one would be \[\frac{1}{5e^5}(2.5-e)^5\]

OpenStudy (anonymous):

as long as this is close enough to zero, you are fine

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=1%2F%285e^5%29%282.5-e%29^5 looks ok

OpenStudy (anonymous):

i guess so. Thanks for your help. You wouldnt be familair with maple would you?